OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation 4x - 3y = 24. Choices... x intercept = (8, 0); y intercept = (0, 6) x intercept = (-8, 0); y intercept = (0, -6) x intercept = (-6, 0); y intercept = (0, 8) x intercept = (6, 0); y intercept = (0, -8) WHICH ONE?!

OpenStudy (anonymous):

well the y intercept is the b when it is in y = mx+b format so subtract 4x on both sides then divide by -3 on both sides

OpenStudy (anonymous):

so you should get y = 4/3x-7 or something like that i might now be tots accurate

OpenStudy (anonymous):

sorry its 8

OpenStudy (anonymous):

y=4/3-8

OpenStudy (precal):

if x = 0 what is y=? that is your y intercept if y=0 what is x=? that is your x intercept

OpenStudy (anonymous):

so then just plug in 0 for y and then that is your x intercept

OpenStudy (anonymous):

thats actually an easier way of doing it but its whatever

OpenStudy (anonymous):

x intercept = (6, 0); y intercept = (0, -8)

OpenStudy (sasogeek):

http://openstudy.com/study#/updates/4f25f23ce4b0a2a9c2673515 <<<< the way to get the answer was explained here by me lol, this question is a duplicate